Bob and Kevin both have square vegetable gardens. Kevin's garden is 10 feet longer than Bob's garden, and its area is four times the area of Bob's garden. What are the dimensions of Bob's garden?
Answer by mananth(16949)

Bob's garden -side length be x feet
Area = x*x=x^2
..
Kevin's garden - side length = x+10 feet
..
so area = (x+10)*(x+10)
=x^+20x+100
..
x^2+20x+100=4x^2
3x^2-20x-100=0
3x^2-30x+10x-100=0
3x(x-10)+10(x-10)=0
(x-10)(3x+10)=0
x= 10
Bob's garden dimensions = 10 feet by 10 feet
...
m.ananth@hotmail.ca